A thin uniform square lamina of side "a" is placed  with its sides parallel to x
and y axis and its center coinciding with the origin.Its moment of inertia about an
axis passing through a point on the y axis at a distance y=2a and parallel to x-axis
is equal to its moment of inertia about an axis passing through a point on the x-axis
at a distance x=d and perpendicular to xy plane . the value of 'd' is

1) 7a/3
2)9a/5
3) Sqrt(47/12)a
4) Sqrt (51/12)a

Grade:12

1 Answers

Uzma kousar
6 years ago
Option 3 : square root (47/12)aFrom parallel axis theorm I=Ig+MR×RFrom perpendicular axis theorm Iz=Ix+Iy